A planet of mass 1.000 × 1026 kg, has a moon of mass 2.100 × 1022 kg and orbital radius 3.550 × 105 km. The orbit of the moon is assumed to be circular.

Determine the distance, from the centre of the planet, the point at which the resultant gravitational field is zero.


1
Expert's answer
2021-03-21T11:31:32-0400
"G\frac{M}{x^2}=G\frac{m}{(R-x)^2}\\\frac{10^4}{x^2}=\frac{2.1}{(3.55-x)^2}\\x=3.499\cdot10^5\ km"